OpenStudy (anonymous):

Determine whether the following sequence converge or diverge. Find the limit if it converges an= n ln(n-2/n) So I think the limit goes to infinity. This would mean that the sequence is divergent, right?

OpenStudy (anonymous):

\[an= n \ln (n-2/n)\]

OpenStudy (anonymous):

Is it n-(2/n) or (n-2)/n?

OpenStudy (anonymous):

(n-2)/n

OpenStudy (anonymous):

So I did \[\lim_{n \rightarrow \infty }n \times \ln ( \lim_{n \rightarrow \infty}(n-2)/n)\]

OpenStudy (anonymous):

I found that the limit on the right hand side is 1, and the left goes to infinity

OpenStudy (dumbcow):

lim((n-2)/n) = 1 ln(1) = 0

OpenStudy (anonymous):

but isn't infinity times 0 undefined?

OpenStudy (dumbcow):

im not sure, i believe it would go to zero

OpenStudy (anonymous):

No, 0 * infinity is indeterminate. Which basically means you have to find a way to solve the limit without falling into that "indetermination"

OpenStudy (anonymous):

hmm, I see

OpenStudy (dumbcow):

you can use L'Hopitals rule first write it as \[\lim_{? \rightarrow ?} \frac{\ln(\frac{n-2}{n})}{\frac{1}{n}}\] Then differentiate top and bottom Then re-evaluate the limit

OpenStudy (anonymous):

goes to 0?

OpenStudy (dumbcow):

No goes to -2 if i did my math right

OpenStudy (anonymous):

dumbcow is right, it goes to -2.

OpenStudy (anonymous):

is the natural log part a common limit?